# Causality-enhanced Decision-Making for Autonomous Mobile Robots in Dynamic Environments

### v3

@@ 5 Experiments
Figure 7: Experimental scenario simulated using our simulator, as described in Sec. 4. The scenario is divided into 11 time slots ($S$), during which worker congestion areas and robot tasks change dynamically.
This section demonstrates how our causality-based decision-making framework (Sec. 3) enhances the robot’s efficiency and safety. In particular, our approach enables the robot to estimate both congestion levels in specific areas of the environment at different times and the battery consumption required for a given task. Estimating congestion allows the robot to anticipate the habitual presence of workers in certain areas at specific times, enabling it to proactively avoid these areas, accommodate human activities, and maintain a safer environment. Additionally, estimating battery consumption helps the robot decide whether to proceed with or abort a task, ensuring it does not risk depleting its battery, failing to complete the assigned task, or being unable to autonomously return to the charging station.
@@ 5.1 Experimental Setup
Collecting real-world data in a warehouse-like environment with multiple workers and dynamic contextual factors would require significant time and financial resources. Given these constraints, we evaluate our decision-making framework in simulation, similarly to other recent works on multi-agent systems and human-robot social navigation [59, 60]. Our simulator enables the creation of different controlled and repeatable HRSI scenarios, allowing for a systematic analysis of how environmental conditions, agent behaviours, and task constraints influence HRSIs. The data generated by multiple simulated scenarios is processed by our framework to learn a causal model. The latter is finally used to perform actual inference and decision-making, as explained in the following sections.
@@
We designed 11 different scenarios with varying worker congestion areas and robot tasks, as illustrated in Fig. 7. These scenarios correspond to 11 different time slots ($S1$-$S11$) throughout the day, capturing realistic variations in worker activity. At each time slot, we assigned a higher probability for waypoints ($W\!s$) to be occupied by workers. This models the fact that, at different times of the day, workers perform specific tasks in designated areas. For instance, in $S1$$S5$ and $S7$$S10$, workers are engaged in normal duties, primarily occupying working areas such as shelves and offices. The slot $S6$ represents lunchtime, during which workers are concentrated in the canteen area. Finally, $S11$ models the off-time period when workers are typically absent. Some congestion areas (represented by a red crowd icon in Fig. 7) correspond to waypoints with the highest probability of being occupied. However, workers are still allowed to move to other locations with lower probabilities, i.e., non-congested areas represented by white circles in Fig. 7. This models the natural behaviour of workers taking breaks, going to the toilet, visiting the canteen, or occasionally moving to other locations beyond their primary working area.
Also the robot’s tasks vary depending on the time slot to align with human activities. In $S1$$S10$ (except $S6$), the robot continuously moves between designated targets (represented as green crossed circles in Fig. 7) at the top and bottom of the working area. Each time, it randomly selects one target from the top and one from the bottom, simulating a pick-and-place task. In $S6$ (lunchtime), the robot assists in the canteen by performing service-related tasks, such as carrying trays or collecting and disposing of waste in bins. The robot’s goals in Fig. 7 ($S6$) reflect this scenario: one near the canteen and the other near the entrance. Finally, in $S11$, when workers are no longer in the warehouse, the robot is assigned a cleaning task. During this period, it navigates through all the waypoints, covering the entire workspace. This final time slot serves to validate the behaviour of our decision-making approach in an empty environment, ensuring that it performs identically to the baseline when no human-robot interactions occur.
@@
We collected a full-day simulated data, including robot’s velocity $V$, battery level $B$, and number of people at each waypoint $W$. Additionally, contextual factors such as the presence of unexpected obstacles $O$ (modelled by red boxes as shown in Fig. 12 top), charging status $C$, waypoint $W$, and time slot $S$ were directly extracted from the simulator. We simulated the presence of 50 workers in the warehouse-like scenario, ensuring that their goal selection aligned with the scenario shown in Fig. 7.
@@
We assumed the robot starts with a fully-charged battery and set its parameters to be consistent with its actual lifetime7 in both idle and motion modes. During this data collection phase, the robot navigated the environment using a ROS-based navigation stack that relies on the A* algorithm as global planner [61, 62], employing the shortest path as its heuristic to reach its goal. This shortest-path strategy serves as a baseline for evaluating our causal decision-making approach, which will be analysed in the following sections.
@@
The data post-processed in the previous step is then used for causal discovery. We assume an underlying first-order Markov process, and therefore consider a one-step time lag in the causal discovery stage, using J-PCMCI+ to handle context variables.
@@
The reconstructed causal model, shown in Fig. 8, is consistent with the expected causal relationships that describe the full scenario explained in Sec. 4.2 and illustrated in Fig. 5. As anticipated, the causal model consists of two subgraphs: one modelling battery dynamics (left part of the figure) and another capturing people density patterns (right part of the figure). The causal discovery algorithm identified also an additional temporal dependency for the people density ${D_{t-1}\rightarrow D_{t}}$ (i.e., an autocorrelation). Although not originally planned, this result is plausible, since congestion levels typically persist over time due to gradual crowd movements. The causal model is then combined with the discretised data to learn the parameters necessary for the inference engine.
@@
Figure 8: Causal model retrieved from the full scenario. Context variables are shown in grey, while system variables are highlighted in orange. The thickness of the arrows and node borders represents the strength of cross- and auto-causal dependencies, respectively, while the numbers on each node and link indicate the time lag of the dependency.
@@ 5.2 Decision-Making Evaluation Setup
Once built, the model is integrated into the causal inference pipeline of our robot decision-making framework. As explained in Sec. 3.3, the inference pipeline estimates key quantities that guide path selection. Specifically, it employs the heuristic defined in Eq. (9) to determine a path that balances distance, congestion avoidance, and battery efficiency. Additionally, it enables the robot to decide whether to execute or abort a task based on its battery level. Our causality-based approach is compared against a common shortest-path heuristic for robot navigation [63, 64], which serves as a benchmark.
@@
An experiment is conducted for each time slot $S$. Workers move according to the waypoint probability distribution defined for $S$, as shown in Fig. 7. Similarly, the robot performs a predetermined number of tasks, always following the plan outlined in the figure. Regardless of the $S$ duration, each experiment concludes when the robot completes all assigned tasks for that time slot, so that baseline and causal approaches are compared on the same number of tasks rather than how many tasks the robot manages to complete within $S$. To eliminate potential biases, each experiment begins with the robot’s battery fully charged. Additionally, we predefine the workers’ goal sequences to match the scenarios in Fig. 7, ensuring that the only difference between the two approaches is the heuristic used for path planning. We set a $45$ seconds deadline for task completion: if the robot fails to complete a task within this time limit, it is considered a failure, and the robot is moved to the next task’s starting position. This prevents the robot from being indefinitely stuck due to nearby human presence, while also ensuring that each task starts from the same location for both approaches.
@@
From $S1$ to $S10$, the number of tasks is set to 200 in order to collect enough interaction data for assessing the benefits of $D$ and battery consumption estimations, as well as evaluating the effect of causal decision-making process. In $S11$, when there are no workers, the robot navigates through all the waypoints in the environment, simulating a simple cleaning routine during non-working hours. Each connection between waypoints is considered as a task that the robot must complete. The minimum number of arcs connecting all the waypoints was determined using the Traveling Salesman Problem [65], resulting in 90 connections, corresponding to 90 tasks.
@@
The heuristic uses the following weights: ${\lambda_{D}=1}$, ${\lambda_{D}=10}$, and ${\lambda_{L}=5}$. These values are empirically chosen to avoid congested areas instead of minimising the total travelled distance, so that the robot favors safer and less crowded paths. If the battery drops below a critical threshold (i.e., $20\%$ of the total charge) during the execution of a task, this is considered failed, and the battery is reset to $100\%$. If the battery falls below the critical threshold after a task has completed, it is still reset but considered a success. Note that for the causal approach, if our decision making framework does not return any valid path to a given goal that satisfies the critical battery constraint, the battery level is set again to $100\%$, as if the robot had been at the charging station.
@@ 5.3 Metrics
Besides principles for social navigation, [66] specifies metrics for their assessment, some of which are used to evaluate the efficiency of our approach:
@@
  • Collision Count: measures the number of robot-human collisions, considered as such when the distance between the robot and a person falls below a predefined safety threshold [67]. Specifically, this metric counts instances where a worker enters the robot’s circumscribed region [^10] http://wiki.ros.org/costmap_2d.
  • Space Compliance: evaluates the robot’s adherence to human spatial preferences based on Hall’s proxemic zones [68].
To assess the significance of the differences between our causal approach and the baseline, we apply opportune statistical tests based on the nature of each metric [69]. In particular, for the Success and Failure Rate, we use a Chi-Square test, which is appropriate for categorical outcomes (success vs. failure). For the Collision Count, we employ a Negative Binomial test, suitable for count data that exhibits overdispersion (i.e., variance greater than the mean), as observed in our case. For all the other continuous or ordinal metrics—Task Time, Travelled Distance, Battery Usage, and Space Compliance—we use the Mann-Whitney U test, a non-parametric alternative to the t-test, since it does not assume normally distributed data.
@@ 5.4 Results
We present the results of our experiments, highlighting the benefits of our approach in in terms of efficiency and safety.
@@
As outlined in Sec. 5.3, the efficiency results shown in Fig. 9 include the following metrics: (top-left) task success and failure counts, (top-right) travelled distance, (bottom-left) task completion time, and (bottom-right) battery consumption. Each graph contains two stacked bars, one for the baseline and one for the causal approach. Each bar is divided into multiple blocks, representing different categories within the same metric. Within each block, the percentage relative to the total height of that bar. This provides a clear visual representation of the proportion of each category within the same bar.
@@
Figure 9: Summary of overall efficiency metrics across all time slots. (Top-left) Task outcomes: (blue) successful tasks, (orange) failures due to high people density $D$, and (red) failures caused by excessive battery consumption $L$. (Top-right) Travelled distance breakdown: (blue) planned distance (expected travel), (orange) extra distance (deviations from the expected), and (red) wasted distance linked to failed tasks. (Bottom-left) Task completion time distribution: (blue) active time (robot moving during successful tasks), (orange) stalled time (robot idle in successful tasks), and (red) wasted time (time spent in failed tasks). (Bottom-right) Battery usage analysis: (blue) effective battery consumption in successful tasks and (red) wasted battery in failed tasks.
@@
Starting with number of successful and failed tasks (top-left), we distinguish the latter into failures due to the robot getting stuck because of people congestion (orange) and failures caused by the battery level dropping below the critical threshold (red).
@@
The travelled distance (top-right) is also divided into three components: the total planned path length across all time slots (blue), representing the total length of the paths determined by the heuristic; extra travelled distance (orange), which quantifies deviations from the planned paths, computed as the difference between the actual total travelled distance and the total planned distance; and wasted distance (red), corresponding to the distance covered during tasks that ultimately failed. The planned and extra travelled distances are only measured for successfully completed tasks. This metric is shown in kilometers (km).
@@
The task time (bottom-left) is divided into three components: active time (blue), representing the total duration in which the robot is actively moving toward its goal (i.e., performing tasks); stalled time (orange), accounting for periods when the robot stopped due to congestion or obstacle avoidance; and wasted time (red), representing the total time spent on tasks that ultimately resulted in a failure. Note that active and stalled times are only measured for successfully completed tasks. This metric is measured in hours (h).
@@
The battery usage (bottom-right) is divided into two categories: effective battery usage (blue), representing the battery consumed during successfully completed tasks, and wasted battery usage (red), which quantifies the battery consumed in tasks that ultimately resulted in a failure. Battery usage is measured in battery cycles, where one cycle corresponds to consuming $100\%$ of the battery.
@@
Sec. 5.3 defines the metrics to evaluate the safety of HRSI, shown in Fig. 10, and including (left) number of collisions and (right) human-robot proxemic compliance. The number of collisions quantifies instances where robot and human are within contact distance, which is TIAGo’s base radius7. The graph compares the baseline (blue) and causal (orange) approaches. The right part of Fig. 10 shows a box plot of the distribution of human-robot distances across all time slots. The background highlights different proxemic zones [68]: public (${\leq 7.6\text{m}}$, green), social (${\leq 3.6\text{m}}$, blue), personal (${\leq 1.2\text{m}}$, yellow), and intimate (${\leq 0.5\text{m}}$, red).

### v4

@@ 5 Experiments
Figure 7: Experimental scenario simulated using our simulator, as described in Sec. 4. The scenario is divided into 11 time-slots ($S$), during which worker congestion areas and robot tasks change dynamically.
This section evaluates the proposed causality-based decision-making framework (Sec. 3) to demonstrate its benefits for robot efficiency and safety in a dynamic, shared environment.
Our framework’s core contribution is twofold: (i) it uses a causal model to improve path selection by scoring routes based on expected congestion and battery drain, and (ii) preemptively refuse tasks that are likely to fail due to low battery. To precisely quantify the individual and combined impact of these two mechanisms, our evaluation is structured as an ablation study. We compare the performance of four distinct approaches:
  • Baseline (Shortest-Path): a standard A* planner using only Euclidean distance, which must accept all tasks;
  • Causal Routing: our causal heuristic (Eq. 9) for path selection, but forced to accept all tasks;
  • Refusal-Only: the baseline planner, but with our causal task refusal mechanism enabled;
  • Full Causal Framework: the complete system, using both causal routing and causal refusal.
This comparative analysis, conducted within our PeopleFlow simulator, clearly isolate the gains from smarter routing versus intelligent task refusal, validating the overall contribution to safety and efficiency.
@@ 5.1 Experimental Setup
All evaluation experiments were conducted on a desktop computer (12th Gen Intel Core i9-12900F × 24, NVIDIA GeForce RTX 4070, 32 GB RAM), based on Ubuntu 20.04 LTS and ROS-Noetic.
Collecting real-world data in a warehouse-like environment with multiple workers and dynamic contextual factors would require significant time and financial resources. Given these constraints, we evaluate our decision-making framework in simulation, similarly to other recent works on multi-agent systems and human-robot social navigation (Mavrogiannis and Knepper, 2019, 2021). Our simulator enables the creation of different controlled and repeatable HRSI scenarios, allowing for a systematic analysis of how environmental conditions, agent behaviours, and task constraints influence HRSIs. The data generated by multiple simulated scenarios is processed by our framework to learn a causal model. The latter is finally used to perform actual inference and decision-making, as explained in the following sections.
@@
We designed 11 different scenarios with varying worker congestion areas and robot tasks, as illustrated in Fig. 7. These scenarios correspond to 11 different time-slots ($S1$-$S11$) throughout the day, capturing realistic variations in worker activity. At each time-slot, we assigned a higher probability for waypoints (WsW\!s) to be occupied by workers. This models the fact that, at different times of the day, workers perform specific tasks in designated areas. For instance, in S1S1S5S5 and S7S7S10S10, workers are engaged in normal duties, primarily occupying working areas such as shelves and offices. The slot $S6$ represents lunchtime, during which workers are concentrated in the canteen area. Finally, $S11$ models the off-time period when workers are typically absent. Some congestion areas (represented by a red crowd icon in Fig. 7) correspond to waypoints with the highest probability of being occupied. However, workers are still allowed to move to other locations with lower probabilities, i.e., non-congested areas represented by white circles in Fig. 7. This models the natural behaviour of workers taking breaks, going to the toilet, visiting the canteen, or occasionally moving to other locations beyond their primary working area.
Also the robot’s tasks vary depending on the time-slot to align with human activities. In S1S1S10S10 (except $S6$), the robot continuously moves between designated targets (represented as green crossed circles in Fig. 7) at the top and bottom of the working area. Each time, it randomly selects one target from the top and one from the bottom, simulating a pick-and-place task. In $S6$ (lunchtime), the robot assists in the canteen by performing service-related tasks, such as carrying trays or collecting and disposing of waste in bins. The robot’s goals in Fig. 7 ($S6$) reflect this scenario: one near the canteen and the other near the entrance. Finally, in $S11$, when workers are no longer in the warehouse, the robot is assigned a cleaning task. During this period, it navigates through all the waypoints, covering the entire workspace. This final time-slot serves to validate the behaviour of our decision-making approach in an empty environment, ensuring that it performs identically to the baseline when no human-robot interactions occur.
@@
Figure 8: Gazebo view of our simulated warehouse-like scenario staged by PeopleFlow. The scenario shows workers (red cylinders) performing tasks in the working area, as well as in the offices, corridor, and canteen. The TIAGo robot navigates between target positions to complete tasks while avoiding unexpected obstacles (purple cubes) in the environment.
We collected a full-day simulated data, including robot’s velocity $V$, battery level $B$, and number of people at each waypoint $W$. Additionally, contextual factors such as the presence of unexpected obstacles $O$ (modelled by purple boxes as shown in Fig. 8), charging status $C$, waypoint $W$, and time-slot $S$ were directly extracted from the simulator. We simulated the presence of 50 workers in the warehouse-like scenario, ensuring that their goal selection aligned with the scenario shown in Fig. 7.
@@
We assumed the robot starts with a fully-charged battery and set its parameters to be consistent with its actual lifetime7 in both idle and motion modes. During this data collection phase, the robot navigated the environment using a ROS-based navigation stack that relies on the A* algorithm as global planner (Marder-Eppstein et al., 2010; Zheng et al., 2023), employing the shortest path as its heuristic to reach its goal. This shortest-path strategy serves as a baseline for evaluating our causal decision-making approach, which will be analysed in the following sections.
@@
The data post-processed in the previous step is then used for causal discovery. We assume an underlying first-order Markov process, and therefore consider a one-step time lag in the causal discovery stage. This Markovian assumption is widely used in the literature for modelling and predicting human spatial behaviours and human-robot collaboration (Asahara et al., 2011; Liu and Wang, 2017; Wakim et al., 2004). We use the causal discovery algorithm J-PCMCI+ to handle context variables. As stated in Sec. 3.2, J-PCMCI+ is specifically designed to model context variables (like $S$, $W$, $C$, $O$) as external factors that affect the system (like $D$, $V$, $L$) without being influenced by it. This structural constraint inherently enforces “expected” real-world directionality by only permitting causal links from context-to-system and system-to-system, while disallowing links from system-to-context (e.g., a time-slot can cause cafeteria congestion, but cafeteria congestion cannot change the time-slot). Specifically, for the conditional independence tests of J-PCMCI+, we employ the RegressionCI test (from the tigramite package [^12] https://github.com/jakobrunge/tigramite) given its ability to handle a mixed dataset of continuous (e.g., velocity) and discrete/categorical (e.g., time-slot) variables, as in our case. We use the standard significance threshold of $\alpha=0.05$ for these tests.
The data post-processed in the previous step is then used for causal discovery. We assume an underlying first-order Markov process, and therefore consider a one-step time lag in the causal discovery stage. This Markovian assumption is widely used in the literature for modelling and predicting human spatial behaviours and human-robot collaboration (Wakim et al., 2004; Asahara et al., 2011; Liu and Wang, 2017). We use J-PCMCI+ to handle context variables. As stated in Sec. 3.2, J-PCMCI+ is specifically designed to model context variables (like $S$, $W$, $C$, $O$) as external factors that influence the system (like $D$, $V$, $L$) without being influenced by it. This structural constraint inherently enforces “obvious” real-world directionality by only permitting causal links from context-to-system and system-to-system, while disallowing links from system-to-context (e.g., a time-slot can cause cafeteria congestion, but cafeteria congestion cannot change the time-slot). Specifically, for the conditional independence tests, we employ the RegressionCI test (from the tigramite package [^13] https://github.com/jakobrunge/tigramite) given its ability to handle a mixed dataset of continuous (e.g., velocity) and discrete/categorical (e.g., time-slot) variables, as in our case. We use the standard significance threshold of $\alpha=0.05$ for these tests.
Note that the J-PCMCI+ algorithm is executed in a single run on the full simulated dataset, as analysing the entire working day was essential to capture all contextual dynamics. This aspect would be missed if smaller data windows were selected for the causal discovery. A detailed consistency analysis of the causal model across different data windows is outside the scope of this paper, whose focus is on the novel end-to-end decision-making framework. However, a detailed analysis of the causal discovery for various data parameters—i.e., time-series length and sampling frequency—was presented in our prior work on ROS-Causal (Castri et al., 2024a).
In terms of computational cost, it is well-established that causal discovery scales with the number of variables (Runge, 2018; Castri et al., 2023, 2024c). For our specific model, which has 7 variables, the J-PCMCI+ algorithm took 4 minutes and 12 seconds to run on the hardware specified in Sec. 5.
@@
The reconstructed causal model, shown in Fig. 9, is consistent with the expected causal relationships that describe the full scenario explained in Sec. 4.2 and illustrated in Fig. 5. As anticipated, the causal model consists of two sub-graphs: one modelling battery dynamics (left part of the figure) and another capturing people density patterns (right part of the figure). The causal discovery algorithm identified also an additional temporal dependency for the people density $D_{t-1}\rightarrow D_{t}$} (i.e., an autocorrelation). Although not originally planned, this result is plausible, since congestion levels typically persist over time due to gradual crowd movements. The causal model is then combined with the discretised data to learn the parameters necessary for the inference engine. As already specified in Sec. 3.3, these steps of causal discovery and parameters learning are performed only once and are not part of the following decision-making process.
In terms of computational cost, it is well-established that causal discovery scales with the number of variables (Runge, 2018; Castri et al., 2023, 2024c). For our specific model, which has 7 variables, the J-PCMCI+ algorithm took 4 minutes and 12 seconds to run on the hardware specified in Sec. 5.
@@
Figure 9: Causal model retrieved from the full scenario. Context variables are shown in grey, while system variables are highlighted in orange. The thickness of the arrows and node borders represents the strength of cross- and auto-causal dependencies, respectively, while the numbers on each node and link indicate the time lag of the dependency. To improve readability, arrows without a number represent a contemporaneous effect (a time lag of 0).
@@ 5.2 Decision-Making Evaluation Setup
As introduced in Sec. 5, our evaluation compares four approaches. The setup for these is as follows. The causal model, shown in Fig. 9, is integrated into the causal inference pipeline of our robot decision-making framework (Fig. 2). As explained in Sec. 3.3 and 4.3, this pipeline estimates key quantities that guide path selection: battery consumption ($\hat{L}$) using Eq. (6) and people density ($\hat{D}$) using Eq. (7). These estimates are not simply reactive predictions: Eq. (6) estimates battery consumption adjusting for the confounder $O$, while Eq. (7) quantifies people density based on time and location.
As a baseline, we benchmark on a common shortest-path heuristic for robot navigation (Nearchou, 1998; Sariff and Buniyamin, 2006; Tan et al., 2021). This approach does not use the causal estimates and is not permitted to refuse tasks. This choice of baseline, while simple, is a standard practice in robotics research (Mombelli et al., 2022; Yang et al., 2020; Apoorva et al., 2018; Ng et al., 2020; Zhang et al., 2024) and reflects the default behaviour of widely-used robotics frameworks like ROS, making the comparison relevant to real-world applications.
Our three causal-aware approaches use the estimates ($\hat{L}$ and $\hat{D}$) to enhance this baseline:
  • Causal Routing: uses the full causal heuristic (Eq. 9) but is forced to accept all tasks (i.e., the refusal mechanism in Eq. (10) is disabled).
  • Refusal-Only: uses the causal task refusal mechanism (Eq. 10) to abort tasks but relies on the simple shortest-path heuristic for routing.
  • Full Causal Framework: employs both the causal heuristic (Eq. 9) and the causal task refusal mechanism (Eq. 10).
@@
An experiment is conducted for each time-slot $S$. Workers move according to the waypoint probability distribution defined for $S$, as shown in Fig. 7. Similarly, the robot performs a predetermined number of tasks, always following the plan outlined in the figure. Regardless of the $S$ duration, each experiment concludes when the robot completes all assigned tasks for that time-slot, so that baseline and causal approaches are compared on the same number of tasks rather than how many tasks the robot manages to complete within $S$. To eliminate potential biases, each experiment begins with the robot’s battery fully charged. Additionally, we predefine the workers’ goal sequences to match the scenarios in Fig. 7, ensuring that the only difference between the two approaches is the heuristic used for path planning. We set a $45$ seconds deadline for task completion: if the robot fails to complete a task within this time limit, it is considered a failure, and the robot is moved to the next task’s starting position. This prevents the robot from being indefinitely stuck due to nearby human presence, while also ensuring that each task starts from the same location for both approaches.
@@
From $S1$ to $S10$, the number of tasks is set to 200 in order to collect enough interaction data for assessing the benefits of $D$ and battery consumption estimations, as well as evaluating the effect of causal decision-making process. In $S11$, when there are no workers, the robot navigates through all the waypoints in the environment, simulating a simple cleaning routine during non-working hours. Each connection between waypoints is considered as a task that the robot must complete. The minimum number of arcs connecting all the waypoints was determined using the Traveling Salesman Problem (Gutin and Punnen, 2006), resulting in 90 connections, corresponding to 90 tasks.
@@
Following a comprehensive sensitivity analysis (presented in B), we selected the heuristic weights: $\lambda_{\delta}=1$}, $\lambda_{D}=10$}, and $\lambda_{L}=5$}. These values were chosen to ensure the robot favours safer and less crowded paths, prioritising congestion avoidance over minimising the total travelled distance. If the battery drops below a critical threshold (i.e., $20\%$ of the total charge) during the execution of a task, this is considered failed, and the battery is reset to $100\%$. If the battery falls below the critical threshold after a task has completed, it is still reset but considered a success. Note that for the causal approach, if our decision making framework does not return any valid path to a given goal that satisfies the critical battery constraint, the battery level is set again to $100\%$, as if the robot had been at the charging station.
@@ 5.3 Metrics
Besides principles for social navigation, (Francis et al., 2024) specifies metrics for their assessment, some of which are used to evaluate the efficiency of our approach:
@@
  • Collision Count: measures the number of robot-human collisions, considered as such when the distance between the robot and a person falls below a predefined safety threshold (Mavrogiannis et al., 2023). Specifically, this metric counts instances where a worker enters the robot’s circumscribed region [^14] http://wiki.ros.org/costmap_2d.
  • Space Compliance: evaluates the robot’s adherence to human spatial preferences based on Hall’s proxemic zones (Hall et al., 1968).
To assess the significance of the differences between our causal approach and the baseline, we apply opportune statistical tests based on the nature of each metric (Rabe-Hesketh and Everitt, 2003). In particular, for the Success and Failure Rate, we use a Chi-Square test, which is appropriate for categorical outcomes (success vs. failure). For the Collision Count, we employ a Negative Binomial test, suitable for count data that exhibits overdispersion (i.e., variance greater than the mean), as observed in our case. For all the other continuous or ordinal metrics—Task Time, Travelled Distance, Battery Usage, and Space Compliance—we use the Mann-Whitney U test, a non-parametric alternative to the t-test, since it does not assume normally distributed data.
@@ 5.4 Results
We present the results of our experiments, highlighting the benefits of our approach in terms of efficiency and safety.
@@
As outlined in Sec. 5.3, the efficiency results shown in Fig. 10 include the following metrics: (10(a)) task success and failure counts, (10(b)) travelled distance, (10(c)) task completion time, and (10(d)) battery consumption. Each graph directly compares the four approaches from our ablation study (Sec. 5.2): the Baseline, Causal Routing, Refusal-Only, and the Full Causal framework. Each approach is represented by a stacked bar, which is divided into multiple blocks representing different categories within that metric. The percentage shown in each block is relative to the total height of its bar, providing a clear visual representation of the proportion of each category.
@@
Figure 10: Summary of overall efficiency metrics across all time-slots. (a) Task outcomes: (blue) successful tasks, (orange) failures due to high people density $D$, and (red) failures caused by excessive battery consumption $L$. (b) Travelled distance breakdown: (blue) planned distance (expected travel), (orange) extra distance (deviations from the expected), and (red) wasted distance linked to failed tasks. (c) Task completion time distribution: (blue) active time (robot moving during successful tasks), (orange) stalled time (robot idle in successful tasks), and (red) wasted time (time spent in failed tasks). (d) Battery usage analysis: (blue) effective battery consumption in successful tasks and (red) wasted battery in failed tasks.
Table 1: Baseline vs Causal runtime performance.
Hardware
CPU: 12th Gen Intel Core i9-12900F
RAM: 32GB
GPU: RTX 4070
OS: Ubuntu 20.04 LTS
ROS: Noetic
Metric Baseline Causal
Avg. Query Time (s) 0 (N/A) 0.327
Avg. Candidate Routes (count) 18.20 25.46
@@
Starting with number of successful and failed tasks (Fig. 10(a)), we distinguish the latter into failures due to the robot getting stuck because of people congestion (orange) and failures caused by the battery level dropping below the critical threshold (red).
@@
The travelled distance (Fig. 10(b)) is also divided into three components: the total planned path length across all time-slots (blue), representing the total length of the paths determined by the heuristic; extra travelled distance (orange), which quantifies deviations from the planned paths, computed as the difference between the actual total travelled distance and the total planned distance; and wasted distance (red), corresponding to the distance covered during tasks that ultimately failed. The planned and extra travelled distances are only measured for successfully completed tasks. This metric is shown in kilometres (km).
@@
The task time (Fig. 10(c)) is divided into three components: active time (blue), representing the total duration in which the robot is actively moving toward its goal (i.e., performing tasks); stalled time (orange), accounting for periods when the robot stopped due to congestion or obstacle avoidance; and wasted time (red), representing the total time spent on tasks that ultimately resulted in a failure. Note that active and stalled times are only measured for successfully completed tasks. This metric is measured in hours (h).
@@
The battery usage (Fig. 10(d)) is divided into two categories: effective battery usage (blue), representing the battery consumed during successfully completed tasks, and wasted battery usage (red), which quantifies the battery consumed in tasks that ultimately resulted in a failure. Battery usage is measured in battery cycles, where one cycle corresponds to consuming $100\%$ of the battery.
In addition to task-level efficiency, we measured the computational cost of the causal query loop, as reported in Table 1. The table summarises the hardware configuration used for the experiments and compares the baseline and causal approaches. In particular, it highlights the runtime performance of the two methods, showing the average query time across all experiments and the average number of candidate routes evaluated by the heuristic functions.
@@
Sec. 5.3 defines the metrics to evaluate the safety of HRSI, shown in Fig. 11, and including (11(a)) number of collisions and (11(b)) human-robot proxemic compliance. The number of collisions quantifies instances where robot and human are within contact distance, which is TIAGo’s base radius7. The graph compares the following approaches: Baseline (red), Causal Routing (blue), Refusal-Only (orange), and Full Causal (green). Fig. 11(b) shows a box plot of the distribution of human-robot distances across all time-slots. The background highlights different proxemic zones (Hall et al., 1968): public ($\leq 7.6\text{m}$}, green), social ($\leq 3.6\text{m}$}, blue), personal ($\leq 1.2\text{m}$}, yellow), and intimate ($\leq 0.5\text{m}$}, red).
Figure 11: Summary of overall safety results across all time-slots. (a) Number of human-robot collisions of the various approaches: Baseline (red), Causal Routing (blue), Refusal-Only (orange), and Full Causal (green). (b) Human-robot proxemic compliance with proxemic zone bands: (green) public, (blue) social, (orange) personal, (red) intimate.